OpenStudy (anonymous):

What is N? n/4 = 8/9 5/8 n = 1/4 2/5 n = 3/5 n /7 = 2/5

OpenStudy (anonymous):

The answer problems are

OpenStudy (anonymous):

A) 2/5 B) 1 1/2 C) 3 5/9 D) 2 4/5

Parth (parthkohli):

n is a number.

Parth (parthkohli):

gotta solve it by multiplying both sides by 36, okay?

Parth (parthkohli):

Or make that step by step: Multiply both by 9. Multiply both by 4.

OpenStudy (anonymous):

ohhh ok so only the bottem parts

Parth (parthkohli):

Bottom parts?

OpenStudy (anonymous):

the bottem part of the fraction 

Parth (parthkohli):

Naw.

Parth (parthkohli):

\[ \color{Black}{\Rightarrow {n \over 4} \times 4 = {8 \over 9} \times 4 }\]Remember that we multiply numerator by numerator and denominator by denominator to multiply two fractions?

OpenStudy (anonymous):

ohhh so multiple the top and bottem by 4

Parth (parthkohli):

Naw.

Parth (parthkohli):

Only the top.

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

ooooo ok i get it lol thx

Parth (parthkohli):

\[ {n \over 4} \times 4 = {n \over 4} \times {4 \over 1} = {n \times 4 \over 4 \times 1} = {4n \over 4} = n \]

Parth (parthkohli):

And same on the right side.

OpenStudy (anonymous):

oh ok thxxx